# A Competition Winning Deep Reinforcement Learning Agent in microRTS

### v1

@@ Appendix A Competition details
Table 5: Policy networks used by RAISocketAI
Network Usage
ppo-Microrts-finetuned-NoWhereToRun-S1-best
NoWhereToRun9x8
ppo-Microrts-A6000-finetuned-coac-mayari-S1-best
All other maps of size 16x16 and smaller
ppo-Microrts-finetuned-DoubleGame-shaped-S1-best
DoubleGame24x24
ppo-Microrts-finetuned-DistantResources-shaped-S1-best
BWDistantResources32x32 if completion time under 75 ms
ppo-Microrts-squnet-DistantResources-128ch-finetuned-S1-best
BWDistantResources32x32 if completion time above 75 ms
ppo-Microrts-squnet-map32-128ch-selfplay-S1-best
All other maps where longest dimension is between 17-32
ppo-Microrts-squnet-map64-64ch-selfplay-S1-best
Maps where the longest dimension is over 32
@@
Table 6: Open competition maps. Representation column is the size of the vectorized observation in RAISocketAI.
Name Size Representation Start
basesWorkers8x8A
8x8
16x16
FourBasesWorkers8x8
8x8
16x16
NoWhereToRun9x8
9x8
12x12* or 16x16
basesWorkers16x16A
16x16
16x16
TwoBasesBarracks16x16
16x16
16x16
DoubleGame24x24
24x24
24x24 or 32x32
BWDistantResources32x32
32x32
32x32
(4)BloodBath.scmB
64x64
64x64
    ppo-Microrts-finetuned-NoWhereToRun-S1-best uses a 12x12 representation
  • ppo-Microrts-A6000-finetuned-coac-mayari-S1-best
  • ppo-Microrts-finetuned-DoubleGame-shaped-S1-best uses a 24x24 representation.
  • ppo-Microrts-squnet-map32-128ch-selfplay-S1-best pads the observation to 32x32.
@@ Appendix B Neural network architecture
DoubleCone(4, 6, 4) ((Limburg, 2023)) consists of (1) 4 residual blocks; (2) a downscaled residual block consisting of a stride-4 convolution, 6 residual blocks, and 2 stride-2 transpose convolutions; (3) 4 residual blocks; and (4) actor and value heads (Figure 1). Each residual block includes a squeeze-excitation layer after the second convolutional layer (Figure 2). The values heads are each (1) 2 stride-2 convolutions, (2) an adaptive average pooling layer, (3) flattened, (4) 2 densely connected layers, and (5) an activation function (Identity [no activation] or Tanh) to a single, scalar value (Figure 3). The adaptive average pooling layer allows the network to be used on different map sizes.
Figure 1: DoubleCone(4, 6, 4) neural network architecture.
Figure 2: ResBlock used in DoubleCone, squnet32, and squnet64. The residual block is similar to a standard residual block but inserts a Squeeze-Excitation block after the convolutional layers and before the residual connection.
Figure 3: Value heads used in (from left to right) DoubleCone, squnet32, and squnet64. The AdaptiveAvgPool2d layer allows the network to be used on other map sizes.
Figure 4: squnet64 neural network architecture. Instead of one downscaling block as in DoubleCone, this network downscales 3 times. This aggressive downscaling reduces the number of computations for larger maps, while theoretically supporting a large receptive field.
Table 7: Comparison of different architectures
DoubleCone squnet-map32 squnet-map64
@@
Deconvolution strides/level [[2, 2]*] [2, 2, 4] [2, 4, 4]
@@ Appendix C Initial training details
Table 8: Initial training schedule from a randomly initialized model
Phase 1 Transition 1→2* Phase 2 Transition 2→3* Phase 3
@@
reward weights [0.8, 0.01, 0.19] [0, 0.5, 0.5] [0, 0.99, 0.01]
$c_{1}$ (value loss coef) [0.5, 0.1, 0.2] [0, 0.4, 0.4] [0, 0.5, 0.1]
@@
Table 9: Comparison of initial training, shaped fine-tuning, and sparse fine-tuning parameters
@@
$\gamma$ (Discount Factor) [0.99, 0.999, 0.999]* " "
GAE $\lambda$ [0.95, 0.99, 0.99] " "
@@
VF Coef Halving True " "
@@
  • Multiply v_loss by 0.5, as done in CleanRL.
@@
Table 10: Shaped fine-tuning schedule
Start Transition →1* Phase 1 Transition 1→2* Phase 2
@@
reward weights [0, 0.99, 0.01] [0, 0.5, 0.5] [0, 0.99, 0.01]
$c_{1}$ (value loss coef) [0, 0.4, 0.2] [0, 0.4, 0.4] [0, 0.5, 0.1]
@@
Table 11: Sparse fine-tuning schedule
Phase 1 Transition 1→2* Phase 2
@@
reward weights [0, 0.99, 0.01] [0, 0.99, 0.01]
$c_{1}$ (value loss coef) [0, 0.5, 0.1] [0, 0.5, 0.1]
$c_{2}$ (entropy coef)) 0.001 0.0001
@@ Appendix D Transfer learning details
Table 12: Transfer learning schedule starting from ppo-Microrts-A6000-finetuned-coac-mayari-S1-best model
Start Transition →1* Phase 1 Transition 1→2* Phase 2
@@
reward weights [0, 0.99, 0.01] [0.4, 0.5, 0.1] [0, 0.99, 0.01]
$c_{1}$ (value loss coef) [0.2, 0.4, 0.2] [0.3, 0.4, 0.1] [0, 0.5, 0.1]
@@
Table 13: Squnet training parameters
@@
Table 14: squnet training schedule starting with randomly initialized weights
Phase 1 Transition 1→2* Phase 2
@@
reward weights [0.8, 0.01, 0.19] [0, 0.99, 0.01]
$c_{1}$ (value loss coef) [0.5, 0.1, 0.2] [0, 0.5, 0.1]
@@ Appendix F Behavior cloning details
Table 15: Neural architecture for behavior cloning and PPO fine-tuned training
@@
Deconvolution strides per level [[2, 2], [2, 2]]*
@@
MACs (16x16) 0.52B
MACs (64x64) 8.40B
@@
Table 16: Behavior cloning training parameters. " means same value as the cell to the left.
@@
Table 17: Training parameters for PPO of behavior cloned models. " means same value as the cell to the left.
@@
VF Coef Halving‡ TRUE " "
@@
Table 18: Behavior cloning schedule for 16x16 maps. Values in transition are linearly interpolated.
@@
Table 19: Behavior cloning schedule for 32x32 and 64x64 maps. Values in transitions are cosine interpolated.
@@
Table 20: Schedule for PPO fine-tuning of behavior cloned model for 16x16 map. Transition values are cosine interpolated.
@@
Table 21: Schedule for PPO fine-tuning of behavior cloned model for 32x32 map. Transition values are cosine interpolated.
@@
Table 22: Schedule for PPO fine-tuning of behavior cloned model for 64x64 map. Transition values are cosine interpolated. Transition 1→2 being empty means values jump from Phase 1 to Phase 2.
@@ Appendix G Training durations
We trained using Lambda Labs GPU on-demand instances. We used single Nvidia GPU instances, but different ones to be able to fit larger minibatches onto the GPU. A10 (24 GB VRAM) and A100 (40 GB VRAM) machines had 30 vCPUs and 200 GB RAM. A6000 (48 GB VRAM) machines had 14 vCPUs and 46 GB RAM. We did not fully utilize the CPU, RAM, or hard drive resources during training.
@@
Table 23: RAISocketAI training durations. Blank models are intermediate models that lead to the next row. For example, the first 3 runs are intermediate models for 16x16. Runs are uploaded to the sgoodfriend/rl-algo-impls-benchmarks Weights and Biases project, except for squnet-DistantResources (sgoodfriend/rl-algo-impls-microrts-2023).
@@
Table 24: Behavior cloning training durations. Runs are uploaded to the sgoodfriend/rl-algo-impls-microrts-2023 Weights and Biases project.
@@
Table 25: Training durations for PPO fine-tuning of behavior cloned models. Runs are uploaded to the sgoodfriend/rl-algo-impls-microrts-2023 Weights and Biases project.
@@ Appendix H Single player round-robin benchmark setup
In Section 4.1, RAISocketAI plays on the 8 Open maps against 4 opponents: (1) baseline POWorkerRush, (2) baseline and 2017 competition winner POLightRush, (3) 2020 competition winner CoacAI, and (4) last competition (2021) winner Mayari. RAISocketAI normally plays against each opponent on each map for 100 matches (50 each as player 1 and 2). The exception is squnet (S) on BWDistantResources32x32, where RAISocketAI only played each opponent for 20 matches (10 each as player 1 and 2). All opponents use A* for pathfinding, which is default for competitions. Win rates are percentages of wins where draws count as 0.5 wins for each player. The single player round-robin benchmark was run on a 2018 Mac Mini with Intel i7-8700B CPU (6-core, 3.2GHz) with PyTorch limited to 6 threads. Timeouts were set to 100 ms. If an agent took 20ms over the deadline (120 ms total), the game was terminated and the win awarded to the opponent.
In Section 4.3, RAI-BC and RAI-BC-PPO play each opponent on each map for 20 matches (10 each as player 1 and 2).
@@ Appendix I Additional IEEE-CoG 2023 microRTS competition details
Table 26: Win rates of all agents in the IEEE-CoG 2023 microRTS competition on Open maps. Player 1 is the row agent and player 2 is the column agent. Each win rate value is the percentage of games won by player 1 (the row agent).
@@
Table 27: RAISocketAI win rates split by competition job. Each map has 5 or 10 jobs that runs a round-robin tournament of all agents for 2 or 1 iterations, respectively. "Outlier" jobs are bolded. "Average" is the average win rate for all jobs. "Average Removing Outliers" is the average win rate for all jobs excluding "outlier" jobs.
basesWorkers8x8A
FourBasesWorkers8x8
NoWhereToRun9x8
basesWorkers16x16A
TwoBasesBarracks16x16
DoubleGame24x24
BWDistantResources32x32
(4)BloodBath.scmB
Overall
98 97 97 100 94 99 49 34
100 94 97 100 4 96 53 38
100 95 33 100 92 100 54 38
5 92 94 98 94 97 58 39
28 95 100 100 94 9 58 19
100 39
97 19
30 44
97 36
95 36
Average 66 95 84 100 75 80 54 34 74
Average Removing Outliers 99 95 97 100 93 98 54 38 84
@@
Table 28: RAI-BC win rate in single player round-robin benchmark. Win rates above 50% are bolded.

### v2

@@ Appendix A Competition details
TABLE VI: Policy networks used by RAISocketAI
Network Usage
ppo-Microrts-finetuned-NoWhereToRun-S1-best NoWhereToRun9x8
ppo-Microrts-A6000-finetuned-coac-mayari-S1-best All other maps of size 16x16 and smaller
ppo-Microrts-finetuned-DoubleGame-shaped-S1-best DoubleGame24x24
ppo-Microrts-finetuned-DistantResources-shaped-S1-best BWDistantResources32x32 if completion time under 75 ms
ppo-Microrts-squnet-DistantResources-128ch-finetuned-S1-best BWDistantResources32x32 if completion time above 75 ms
ppo-Microrts-squnet-map32-128ch-selfplay-S1-best All other maps where longest dimension is between 17-32
ppo-Microrts-squnet-map64-64ch-selfplay-S1-best Maps where the longest dimension is over 32
@@ Appendix B Neural network architecture
DoubleCone(4, 6, 4) [11] consists of (1) 4 residual blocks; (2) a downscaled residual block consisting of a stride-4 convolution, 6 residual blocks, and 2 stride-2 transpose convolutions; (3) 4 residual blocks; and (4) actor and value heads (Supplemental Fig. 1). Each residual block includes a squeeze-excitation layer after the second convolutional layer (Supplemental Fig. 2). The values heads are each (1) 2 stride-2 convolutions, (2) an adaptive average pooling layer, (3) flattened, (4) 2 densely connected layers, and (5) an activation function (Identity [no activation] or Tanh) to a single, scalar value (Supplemental Fig. 3). The adaptive average pooling layer allows the network to be used on different map sizes.
Supplemental Fig. 1: DoubleCone(4, 6, 4) neural network architecture.
Supplemental Fig. 2: ResBlock used in DoubleCone, squnet32, and squnet64. The residual block is similar to a standard residual block but inserts a Squeeze-Excitation block after the convolutional layers and before the residual connection.
Supplemental Fig. 3: Value heads used in (from left to right) DoubleCone, squnet32, and squnet64. The AdaptiveAvgPool2d layer allows the network to be used on various map sizes.
Supplemental Fig. 4: squnet64 neural network architecture. Instead of one downscaling block as in DoubleCone, this network downscales 3 times. This aggressive downscaling reduces the number of computations for larger maps, while theoretically supporting a large receptive field.
Supplemental Table I: Comparison of different architectures
DoubleCone squnet-map32a squnet-map64
@@
Deconvolution strides/level [[2, 2]b ] [2, 2, 4] [2, 4, 4]
@@ Appendix C Initial training details
Supplemental Table II: Initial training schedule from a randomly initialized model
Phase 1 Transition 1→2a Phase 2 Transition 2→3a Phase 3
@@
Reward Weightsb [0.8, 0.01, 0.19] [0, 0.5, 0.5] [0, 0.99, 0.01]
$c_{1}$ (Value Loss Coef)b [0.5, 0.1, 0.2] [0, 0.4, 0.4] [0, 0.5, 0.1]
@@
Supplemental Table III: Comparison of initial training, shaped fine-tuning, and sparse fine-tuning parameters
@@
$\gamma$ (Discount Factor) [0.99, 0.999, 0.999]a " "
GAE $\lambda$ [0.95, 0.99, 0.99]a " "
@@
VF Coef Halvingb True " "
@@
  • b
    Multiply $v_{\text{loss}}$ by 0.5, as done in CleanRL.
  • c
@@
Supplemental Table IV: Shaped fine-tuning schedule
Start Transition →1a Phase 1 Transition 1→2a Phase 2
@@
Reward Weightsb [0, 0.99, 0.01] [0, 0.5, 0.5] [0, 0.99, 0.01]
$c_{1}$ (Value Loss Coef)b [0, 0.4, 0.2] [0, 0.4, 0.4] [0, 0.5, 0.1]
@@
Supplemental Table V: Sparse fine-tuning schedule
Phase 1 Transition 1→2a Phase 2
@@
Reward Weightsb [0, 0.99, 0.01] [0, 0.99, 0.01]
$c_{1}$ (Value Loss Coef)b [0, 0.5, 0.1] [0, 0.5, 0.1]
$c_{2}$ (Entropy Coef) 0.001 0.0001
@@ Appendix D Transfer learning details
Supplemental Table VI: Transfer learning schedule starting from ppo-Microrts-A6000-finetuned-coac-mayari-S1-best model
Start Transition →1a Phase 1 Transition 1→2a Phase 2
@@
Reward Weightsb [0, 0.99, 0.01] [0.4, 0.5, 0.1] [0, 0.99, 0.01]
$c_{1}$ (Value Loss Coef)b [0.2, 0.4, 0.2] [0.3, 0.4, 0.1] [0, 0.5, 0.1]
@@
Supplemental Table VII: Squnet training parameters
@@
Supplemental Table VIII: squnet training schedule starting with randomly initialized weights
Phase 1 Transition 1→2a Phase 2
@@
Reward Weightsb [0.8, 0.01, 0.19] [0, 0.99, 0.01]
$c_{1}$ (Value Loss Coef)b [0.5, 0.1, 0.2] [0, 0.5, 0.1]
@@ Appendix F Behavior cloning details
Supplemental Table IX: Neural architecture for behavior cloning and PPO fine-tuned training
@@
Deconvolution strides per level [[2, 2], [2, 2]]a
@@
MACsb (16x16) 0.52B
MACsb (64x64) 8.40B
  • a
    2 stride-2 transpose convolutions to match each 1 stride-4 convolution.
  • b
@@
Supplemental Table X: Behavior cloning training parameters
@@
    Same value as cell to left.
Supplemental Table XI: Training parameters for PPO fine-tuning of behavior cloned models
@@
VF Coef Halvinga TRUE " "
@@
    Same value as cell to left.
  • a
    Multiply $v_{\text{loss}}$ by 0.5, as done in CleanRL.
Supplemental Table XII: Behavior cloning schedule for 16x16 maps. Values in transition are linearly interpolated.
@@
Supplemental Table XIII: Behavior cloning schedule for 32x32 and 64x64 maps. Values in transitions are cosine interpolated.
@@
Supplemental Table XIV: Schedule for PPO fine-tuning of behavior cloned model for 16x16 map. Transition values are cosine interpolated.
@@
Supplemental Table XV: Schedule for PPO fine-tuning of behavior cloned model for 32x32 map. Transition values are cosine interpolated.
@@
Supplemental Table XVI: Schedule for PPO fine-tuning of behavior cloned model for 64x64 map. Transition values are cosine interpolated. Transition 1→2 being empty means values jump from Phase 1 to Phase 2.
@@ Appendix G Training durations
We trained using Lambda Labs GPU on-demand instances. We used single Nvidia GPU instances, but different GPUs to be able to fit larger minibatches onto the GPU. A10 (24 GB VRAM) and A100 (40 GB VRAM) machines had 30 vCPUs and 200 GB RAM. A6000 (48 GB VRAM) machines had 14 vCPUs and 100 GB RAM. We did not fully utilize the CPU, RAM, GPU compute, or hard drive resources during training.
@@
Supplemental Table XVII: RAISocketAI training durations. Blank models are intermediate models that lead to the next row. For example, the first 3 runs are intermediate models for 16x16. Runs are uploaded to the sgoodfriend/rl-algo-impls-benchmarks Weights and Biases project, except for squnet-DistantResources (sgoodfriend/rl-algo-impls-microrts-2023).
@@
Supplemental Table XVIII: Behavior cloning training durations. Runs are uploaded to the sgoodfriend/rl-algo-impls-microrts-2023 Weights and Biases project.
@@
Supplemental Table XIX: Training durations for PPO fine-tuning of behavior cloned models. Runs are uploaded to the sgoodfriend/rl-algo-impls-microrts-2023 Weights and Biases project.
@@ Appendix H Single player round-robin benchmark setup
In Section IV-A, RAISocketAI plays on the 8 Open maps against 4 opponents: (1) baseline POWorkerRush, (2) baseline and 2017 competition winner POLightRush, (3) 2020 competition winner CoacAI, and (4) last competition (2021) winner Mayari. RAISocketAI normally plays against each opponent on each map for 100 matches (50 each as player 1 and 2). The exception is the squnet model for BWDistantResources32x32, where RAISocketAI only played each opponent for 20 matches (10 each as player 1 and 2). All opponents use A* for pathfinding, which is default for competitions. Win rates are percentages of wins where draws count as 0.5 wins for each player. The single player round-robin benchmark was run on a 2018 Mac Mini with Intel i7-8700B CPU (6-core, 3.2GHz) with PyTorch limited to 6 threads. Timeouts were set to 100 ms. If an agent took 20ms over the deadline (120 ms total), the game was terminated and the win awarded to the opponent.
In Section IV-C, RAI-BC and RAI-BC-PPO play each opponent on each map for 20 matches (10 each as player 1 and 2).
@@ Appendix I Additional IEEE-CoG 2023 microRTS competition details
Supplemental Table XX: Win rates of all agents in the IEEE-CoG 2023 microRTS competition on Open maps. Player 1 is the row agent and player 2 is the column agent. Each win rate value is the percentage of games won by player 1 (the row agent). Higher win rates are redder. Lower win rates are bluer.
@@
Supplemental Table XXI: RAI-BC win rate in single player round-robin benchmark. Win rates above 50% are bolded. Higher win rates are redder. Lower win rates are bluer.
@@
Appendix K Videos of RAISocketAI Against Mayari (2021 CoG Winner)
At https://github.com/sgoodfriend/rl-algo-impls/tree/8230a7c1/papers/cog2024/vsMayariVideos, we provide videos of RAISocketAI playing against the 2021 CoG competition winner Mayari on each of the Open maps. RAISocketAI is always Player 1 in these videos, thus RAISocketAI’s units have a blue outline while Mayari’s units have a red outline. RAISocketAI’s units start at the left or upper part of the map, except in (4)BloodBath.scmB where RAISocketAI starts at the bottom-right. The videos are named after the map they are played on.